what quantity x of a 65% acid solution must mix with a 20% solution to produce 300 mL of a 45% solution
Answer by nerdybill(7384)

what quantity x of a 65% acid solution must mix with a 20% solution to produce 300 mL of a 45% solution
.
Let x = amount (mL) of 65% solution
then
300-x = amount (mL) of 20% solution
.
.65x + .20(300-x) = .45(300)
.65x + 60-.20x = 135
.45x + 60 = 135
.45x = 75
x = 166.67 mL (65% solution)
.
amount of 20% solution:
300-x = 300-166.67 = 133.33 mL